# Electrodynamics of Vortices in Quasi-2D Scalar Bose-Einstein Condensates

### v3

@@ VI Temporal change of the circulation in a static area
VII Non-rotating quasi-2D scalar BEC within the Gross-Pitaevskii equation and the point-vortex model
Given that $\nabla\cdot\bm{v}_{s}\left(\bm{r},t\right)=0$ in the GPE for non-rotating scalar BEC [32], we may introduce using Eq. (24) a differentiable real function $C\left(\bm{r},t\right)$ such that
@@ VII Non-rotating quasi-2D scalar BEC within the Gross-Pitaevskii equation and the point-vortex model
VII.1 Homogeneous BEC number density limit
Let us ignore the spatial fluctuation of $n\left(\bm{r},t\right)$ for simplicity. Of course, this limit is valid only for vanishing healing length $\xi_{h}\rightarrow 0$; else, the spatial fluctuation of $n\left(\bm{r},t\right)$ cannot be neglected since $n\left(\bm{r},t\right)=0$ at cores of vortices. This neglect of the size of the vortex is another assumption in the PVM.
In this limit, $n\left(\bm{r},t\right)$ does not change in time according to Eq. (70), making it to be constant both in space and time. To make $n\left(\bm{r},t\right)$ constant, $V\left(\bm{r},t\right)$ should also be constant. With those conditions, $\bm{H}_{\rm sf}\left(\bm{r},t\right)\simeq 0$ according to the definition in Eq. (44). Also, $\bm{J}_{m,\rm{sf}}\left(\bm{r},t\right)\simeq 0$ and hence $\bm{A}_{m,\rm{sf}}\left(\bm{r},t\right)\simeq 0$, leading to a constant $\bm{A}_{e,\rm{sf}}\left(\bm{r},t\right)$ in space. As $\bm{H}_{\rm sf}\left(\bm{r},t\right)\simeq 0$, the effective Poynting vector is approximately zero, meaning that there is no effective radiation and thus the total vortex number is fixed (individual vortex charge is conserved). It is known that the vortices in non-rotating quasi-2D scalar BEC cannot collide in the PVM, leading to the vortex number conservation in the GPE + PVM description [41, 42]. Our duality reassures these known results.
@@ VII.1 Homogeneous BEC number density limit
VII.2 Relation to 2D Coulomb gas and vortex spacing distribution
As we showed in Eq. (72), the static vortices in the non-rotating quasi-2D BEC in the GPE + PVM description can be mapped to the 2D Coulomb gas. Using the duality between the 2D Coulomb gas and the (1+1) dimensional sine-Gordon model [88, 89], the equivalence between the sine-Gordon model and the massive Thirring model [90], and the work in Ref. [91] that connected between the Berezinskii-Kosterlitz-Thouless (BKT) transition [92, 93] and the 2D Coulomb gas, one can infer that the BKT transition may happen in the non-rotating quasi-2D BEC with vortices. Using the results in Ref. [94], if the topological charge of the vortex is ±$\pm Q$ and the total charge is zero, it can be shown that the BKT transition occurs when the effective temperature $T_{\rm eff}$ of the system is at $T_{c}$, which takes the value
@@ VIII Conclusion
An interesting prospect is the generalization of our results to BEC characterized by higher-order nonlinearities (e.g., due to losses and confinement), dipolar interactions [109], and spinor degrees of freedom [8] where the mean-field wavefunction $\psi\left(\bm{r},t\right)=\left[\begin{array}[]{ccc}\psi_{f}\left(\bm{r},t\right)&\cdots&\psi_{-f}\left(\bm{r},t\right)\end{array}\right]^{T}$ for spin $f$ system does not commute with its hermitian conjugate $\psi^{\dagger}\left(\bm{r},t\right)$ in general. Beyond the realm of ultracold gases, our findings can be applied and generalized for polaritonic BEC and quantum fluids of light [110]. The extending of our results beyond the quasi-2D case is also an open problem. For example, in 3D, one can infer from Eqs. (46) and (47) that there would be effective free magnetic charge since $\bm{e}_{\perp}$ would correspond to the unit vector along the vortex line, which is both space- and time-dependent as the vortex line can be open or closed [66, 12, 111, 112, 113]. This effective free magnetic charge would affect the effective magnetic field strength $\bm{H}_{\rm sf}\left(\bm{r},t\right)$, introducing additional terms to the vortex line interaction in Refs. [32, 33] that neglected vortex core regions.
@@ Acknowledgements.
It is a pleasure to thank Michael Berry, Bogdan Damski, Íñigo L. Egusquiza, Uwe R. Fischer, Étienne Fodor, András Grabarits, Woo Jin Kwon, Matteo Massaro, Kazutaka Takahashi, Mithun Thudiyangal, and Masahito Ueda for insightful comments and discussions. This research was funded by the Luxembourg National Research Fund (FNR), grant reference 17132060. For the purpose of open access, the author has applied a Creative Commons Attribution 4.0 International (CC BY 4.0) license to any Author Accepted Manuscript version arising from this submission.
@@
The 3rd to 5th lines in Eq. (83) can be written as
@@
The 6th to the last lines in Eq. (83) can be written as

### v4

@@ VII Stability of vortices in quasi-2D scalar BEC
VIII Nonrotating quasi-2D scalar BEC within the Gross-Pitaevskii equation and the point-vortex model
Given that $\nabla\cdot\bm{v}_{s}\left(\bm{r},t\right)=0$ in the GPE for nonrotating scalar BEC [34], we may introduce using Eq. (24) a differentiable real function $C\left(\bm{r},t\right)$ such that
@@ VIII Nonrotating quasi-2D scalar BEC within the Gross-Pitaevskii equation and the point-vortex model
VIII.1 Homogeneous BEC number density limit
Let us ignore the spatial fluctuation of $n\left(\bm{r},t\right)$ for simplicity. Of course, this limit is valid only for vanishing healing length $\xi_{h}\rightarrow 0$; else, the spatial fluctuation of $n\left(\bm{r},t\right)$ cannot be neglected since $n\left(\bm{r},t\right)=0$ at cores of vortices. This neglect of the size of the vortex is another assumption in the PVM. As we already discussed in Sec. VII, the effective Poynting vector is approximately zero in this limit, reassuring the known results that the vortices in nonrotating quasi-2D scalar BEC cannot collide in the PVM, thus leading to the vortex number conservation in the GPE + PVM description [43, 44].
@@ VIII.1 Homogeneous BEC number density limit
VIII.2 Relation to 2D Coulomb gas and vortex spacing distribution
As we showed in Eq. (72), the static vortices in the nonrotating quasi-2D BEC in the GPE + PVM description can be mapped to the 2D Coulomb gas. Using the duality between the 2D Coulomb gas and the (1+1) dimensional sine-Gordon model [90, 91], the equivalence between the sine-Gordon model and the massive Thirring model [92], and the work in Ref. [93] that connected between the Berezinskii-Kosterlitz-Thouless (BKT) transition [94, 95] and the 2D Coulomb gas, one can infer that the BKT transition may happen in the nonrotating quasi-2D BEC with vortices. Using the results in Ref. [96], if the topological charge of the vortex is ±$\pm Q$ and the total charge is zero, it can be shown that the BKT transition occurs when the effective temperature $T_{\rm eff}$ of the system is at $T_{c}$, which takes the value
@@ IX Conclusion
An interesting prospect is the generalization of our results to BEC characterized by higher-order nonlinearities (e.g., due to losses and confinement), dipolar interactions [111], and spinor degrees of freedom [8], where the mean-field wavefunction $\psi\left(\bm{r},t\right)=\left[\begin{array}[]{ccc}\psi_{f}\left(\bm{r},t\right)&\cdots&\psi_{-f}\left(\bm{r},t\right)\end{array}\right]^{T}$ for a spin-$f$ system does not commute with its Hermitian conjugate $\psi^{\dagger}\left(\bm{r},t\right)$ in general.
Beyond the realm of ultracold gases, our findings can be applied and generalized for polaritonic BEC and quantum fluids of light [112]. The extending of our results beyond the quasi-2D case is also an open problem. For example, in 3D, one can infer from Eqs. (46) and (47) that there would be effective free magnetic charge since $\bm{e}_{\perp}$ would correspond to the unit vector along the vortex line, which is both space- and time-dependent as the vortex line can be open or closed [68, 14, 113, 114, 115]. This effective free magnetic charge would affect the effective magnetic field strength $\bm{H}_{\rm sf}\left(\bm{r},t\right)$, introducing additional terms to the vortex line interaction in Refs. [34, 35] that neglected vortex core regions.
@@ Acknowledgements.
It is a pleasure to thank Michael V. Berry, Ashton Bradley, Bogdan Damski, Íñigo L. Egusquiza, Uwe R. Fischer, Étienne Fodor, András Grabarits, Woo Jin Kwon, Matteo Massaro, Kazutaka Takahashi, Mithun Thudiyangal, and Masahito Ueda for insightful comments and discussions. The authors acknowledge financial support from the National Research Fund of Luxembourg under Grant No. C22/MS/17132060/BeyondKZM. For the purpose of open access, the author has applied a Creative Commons Attribution 4.0 International (CC BY 4.0) license to any Author Accepted Manuscript version arising from this submission.
@@
The third to fifth lines in Eq. (83) can be written as
@@
The sixth to the last lines in Eq. (83) can be written as